OpenStudy (anonymous):

which order pair is a solution of the equation y=-5x+4? A. (2,14) B. (5,-29) C.(3,-19) D.(5,-21)

OpenStudy (anonymous):

d

OpenStudy (anonymous):

-21=-5(5) + 4 -21= -25 + 4 -21 = - 21

OpenStudy (anonymous):

Thank You!:)